# How structured are the representations in transformer-based vision encoders? An analysis of multi-object representations in vision-language models

### v2

@@ 5 Discussion
In this work, we started by identifying three characteristics required in representations of image models for them to be symbol-like: 1) The architecture should be able to aggregate information specific to objects on the image into smaller units; 2) The architecture should represent various logical units of the input image in separate computational units, i.e., there should be object-wise information disentanglement; 3) A downstream processing network should be able to utilise the composition of these units for the required task.
@@
Most vision encoders fulfill the first criterion we described as they aggregate useful information from the image into smaller representation space useful for the downstream task. However, we were interested in whether separate representational units specialise in representing the image’s separate constituent parts (objects). Our analysis shows that the object-specific (avg_obj, random_obj) tokens show decent decoding accuracy across the models. Each object-wise token’s accuracy for classifying itself is our estimate for fulfilling this criterion. For the second characteristic, we looked at the network’s capability to form disentangled representations of objects. We saw evidence that the information about other objects also leaks into the object-wise representations. Ideally, each object representation can only decode that particular object and other objects in the scene cannot be decoded by that representation. As the other objects can be decoded in our object representations (with accuracies of 0.89 by secondary object representations and 0.65 by primary object representations), their representations are already affected by the context, i.e. other objects. This finding implies that the downstream task using these representations compositionally still cannot perform structured composition as with abstract symbols. Further, the higher the decoding accuracy of other objects in the scene, the higher the entanglement in the representation. The entangled object representation will affect the downstream network’s outputs if the context information is irrelevant to the task.
Our comparative analysis of the BLIP model’s performance showed a notable degradation in decoding efficacy across all representations when objects were omitted from its captions. This observation underscores the significant influence of the model’s downstream objective on its representational capabilities. On the one hand, this enhances the disentanglement of object representations, reducing interference from objects insignificant to the task. Conversely, it reduces the model’s generalizability, particularly when deployed in diverse tasks or faced with unfamiliar input distributions. Our results show that this problem is most exacerbated in the representation of the CLS token. Most research predominantly assesses these networks’ capabilities based on their final outputs, which rely on representations like a CLS token (for example in [28, 15]). Our findings show a loss in representing multiple objects during this process of funneling information into the CLS token. Further, the downstream objective’s effect is the most extreme on the nature of these representations. We show that this causes suboptimal performance of these networks in simple tasks such as retrieval in the multi-object setting. We further postulate that these phenomena (non-binding of certain objects representations or entangled representations) significantly contribute to the models’ suboptimal performance in tasks that require to compose using the representation of objects in scenarios involving multiple objects. In evaluations of CLIP, FLAVA and similar models within a multi-object context, the ramifications of the CLS token bottleneck and its training procedure (training data and objective) must be meticulously considered. The objects represented and the nature of their representation in this bottleneck are heavily influenced by the downstream objective or a co-occurring object, thereby affecting the accuracy of outputs in tasks that diverge from the types of instances encountered during training. Future work can utilise our method to look into training better readouts and fine-tuning on specific underrepresented object images.
@@ 6 Limitations
The probing and representation analysis methods we used in our analysis have some limitations, and the results must be carefully interpreted. High decoding accuracy in a decoding task with two or four objects may be due to the easy decoding task (if the object classes are naturally distinct and it is only a 2/4 way classification). Further, the fact that one can classify a particular object from a few other classes may depend on representing a single or a few features distinguishing between the classes. This representation may not model many other aspects of the object. We controlled for this limitation by training global probes for 20-class classification using the same representations, showing reasonably high accuracy. Likewise, a lower decoding accuracy does not mean that the model does not have information about the particular object; the information is just not encoded with linear distinction at that layer/token. Overall, we are averaging over six tasks in many object categories. Hence, even though the exact numbers may not reflect the model’s exact state, the relative numbers are helpful in making inferences about the model and its representations.
@@ 7 Conclusion
In this work, we first formulated three characteristics for symbolic representations in an image encoder. We then analysed the representations of transformer-based image encoders in VIT, BLIP, CLIP, FLAVA and CLIP (Resnet X 4) image encoders. Through probing tasks, we observed that object-specific areas hold the most discriminative features about the objects until the last layers. Their discriminative ability decreased when the objects were not important for the downstream task but was still reasonably maintained. The token representations are not disentangled since they can decode other objects in the scene with accuracy far above random guesses. We found that the aggregate image representation, i.e. CLS token for transformers, does not represent all objects, and its discriminative ability degrades the most when the object is not useful for the trained downstream task. Our work thus characterises the extent to which the representations of these image encoders are symbolic and gives insights into the failures on out-of-distribution downstream tasks that utilise the composition of these features.

### v3

@@ 5 Discussion
In this work, we identify two properties required in representations of image encoders for them to be more structured: 1) The model should be able to bind information specific to particular objects in the image into specific representation units, i.e. a limited number of tokens can represent the object better than its input representation. 2) The model should segregate various objects of the input image in separate sets of tokens, i.e., there should be object-wise information disentanglement in the token space.
Most vision encoders fulfill the first criterion, as they aggregate information from the image into smaller representation spaces functional for the downstream task. In particular, the decoding accuracy for all objects is higher than that of their raw input representation. However, our results demonstrated that the bottleneck at the CLS token does not represent all objects well.
We were interested in understanding whether separate representational units specialise in representing the image’s constituent parts (objects). Our analysis shows that the object-specific (avg_obj, random_obj) tokens show decent decoding accuracy across the models. Each object-wise token’s accuracy for classifying itself is our estimate for fulfilling this criterion. This accuracy is better than for the CLS token for all models except CLIP-L.
@@
For the second characteristic, we looked at the network’s capability to form disentangled representations of objects in the token space. We saw evidence that information about other objects also leaks into the object-wise representations. Ideally, each object representation can only decode that particular object, and other objects in the scene cannot be decoded by that representation. As the other objects can be decoded in our object representations (with accuracies of 0.89 by secondary object representations and 0.65 by primary object representations), their representations are already affected by the context, i.e. other objects. In addition, the higher the decoding accuracy of other objects in the scene, the higher the entanglement in the representation.
Our comparative analysis of the BLIP model’s performance showed a notable degradation in decoding efficacy across all representations when objects were omitted from its captions. We replicate this result using the object-wise tokens for all models examined in the global probe study. This observation underscores the significant influence of the model’s downstream objective on its representational capabilities. Most research predominantly assesses these networks’ capabilities based on their final outputs, which rely on representations like a CLS token (for example, in [39, 19]). Our findings show a loss in representing multiple objects while funnelling information into the CLS token. Further, the downstream objective’s effect is the most extreme on the nature of the CLS representations.
We show that this causes suboptimal performance of these networks in simple tasks such as identifying objects in the multi-object setting, especially for background objects. We show that the networks have some structured information in token space i.e. they preserve object-specific information localised in the token space. Our analysis finally shows that the measures of structured information in the token space of models correlate with the ability of the models to generalise on background objects using a 20-class experiment. This way, our measures of structured information evaluate the models’ ability to represent individual objects in a multi-object scene. These measures show which model’s representations and at what layer are most structured for representing individual objects separately. They achieve this by not combining object representations in the token space (other than CLS token) under the influence of objective functions that can lead to abstracting the background objects while modelling the image. Access to this structured representation helps with tasks requiring access to detailed representations of individual objects in the scene.
The natural next step of this work is to verify the usefulness of the two representation measures on specific datasets (object categories) for the choice of the best layer and network on actual downstream tasks like semantic segmentation or object localisation. This way, one can estimate if the best-predicted layers for the networks give the best results, or in the case of the use of deep decoders for segmentation like in [37, 20, 4], it will show how much less data is needed to train a decoder as opposed to using the default final layer (especially for good performance on less prominent background objects).
@@ 6 Limitations
The probing and representation analysis methods we used have some limitations, and the results must be carefully interpreted. High decoding accuracy in a decoding task with two or four objects may be due to the easy decoding task (if the object classes are naturally distinct and it is only a 2/4 way classification). Further, the fact that one can classify a particular object from a few other classes may depend on representing a single or a few features distinguishing between the classes. This representation may not model many other aspects of the object. We controlled for this limitation by training global probes for 20-class classification using the same representations, showing reasonably high accuracy. Likewise, a lower decoding accuracy does not mean that the model does not have information about the particular object; the information is just not encoded with linear distinction at that layer/token. Overall, we are averaging over six object sets with many object categories. Hence, even though the exact numbers may not reflect the model’s exact state, we can safely rely on the relative numbers to make inferences about the model and its representations.
@@ 7 Conclusion
In this work, we first formulated two characteristics for structured representations in an image encoder. We then analysed the representations of transformer-based image encoders in VIT, BLIP, CLIP, FLAVA, DINO, DINOv2 and CLIP (Resnet X4) image encoders. Through object decoding tasks, we create a view of the representation structure of these pre-trained models. We observed that object-specific areas hold the most discriminative features about the objects. Their discriminative ability decreased when the objects were not important for the downstream task but was still reasonably maintained. The token representations are not disentangled since they can decode other objects in the scene with accuracy far above random guesses. We found that the aggregate image representation, i.e. CLS token for transformers, does not represent all objects, and its discriminative ability degrades the most when the object is not useful for the trained downstream task. Our work thus characterises and measures the extent to which the representations of these image encoders are structured in the token space, that the individual objects are represented separately in the token space, but their representations are entangled. We also calculated the layerwise measures of representation and entanglement by object importance (estimated using captions), which provides insights into the failures and optimal adaptation of these networks to downstream tasks that utilise these individual object features.
{credits}
7.0.1 Supplementary Materials
The code to obtain data splits from the COCO dataset, experimental setup, and data analysis can be found in the repository linked below. Further detailed plots for the paired object task showing the decoding accuracy of objects across the layers of the network have been given in the notebooks in the supplementary material for all models.
7.0.2 Acknowledgements
We thank Raul Vicente, and Meelis Kull for the valuable discussions and inputs on the paper. This work was supported by the Estonian Research Council grant PSG728, the Estonian Centre of Excellence in Artificial Intelligence (EXAI), funded by the Estonian Ministry of Education and Research and the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreement No. 952060 (Trust AI).
[…]
The authors declare that they have no competing interests.